OpenStudy (anonymous):

find three consecutive positive even integers such that the product of the second and third integera is twenty more than ten times the first integer. Someone please help.me.

OpenStudy (anonymous):

Someone pleaee help.mee

OpenStudy (mertsj):

x= first even integer x+2 = second even integer x+4= third even integer (x+2)(x+4) = product of the second and third integers \[(x+2)(x+4)=10x+20\]

OpenStudy (anonymous):

I got x^2 -4x-12 =O

OpenStudy (mertsj):

\[x^2+6x+8=10x+20\] \[x^2-4x-12=0\] \[(x-6)(x+2)=0\] \[x=6, x+2=8, x+4=10\] \[x=-2, x+2=0, x+4=2\]

OpenStudy (mysesshou):

Good ! last line unused for answer because not all positive even integers

OpenStudy (anonymous):

Im sorry but I dont get it

OpenStudy (mertsj):

You got the right equation. Factor it and solve.

OpenStudy (mertsj):

Can you factor it?

OpenStudy (anonymous):

I only got up to (x-6) (x+2) = O

OpenStudy (mertsj):

If you multiply two numbers and get 0 then one of those numbers has to be 0. So write x-6=0 or x+2=0

OpenStudy (anonymous):

You would get x^2 - 4x - 12 = O

OpenStudy (mysesshou):

x-6=0 x+2=0 solve for x first

OpenStudy (mertsj):

When you solve x-6=0 you get x^2-4x-12=0?????????

OpenStudy (anonymous):

Ur answer wud be x = 6 and x = -2

OpenStudy (mertsj):

Yes. But -2 is not positive so discard it and then find the other two consecutive even integers if the first one is 6